# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ Appendix H Proof of The Main Theorems
Choose a positive integer $m_{0}$ such that, for $10<\gamma<$ 1,
$\iota^{-2m_{0}}\geq(1-\gamma)^{-1}\max\left\{\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)^{-1}\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}+m\right),\frac{L_{gy}^{2}}{\mu_{g}^{2}}\tilde{\kappa}\left(\frac{\mu_{g}L_{fy}+C_{fy}L_{gyy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\right)^{2}\right\},$
which means $m_{0}\sim\mathcal{O}\left(\log m\right)$.
Since we set the step size for $x$ as zero in the first $m_{0}$ steps, 3.10 reveals that
$\left(\varepsilon_{j}^{h}\right)^{2}\leq\lambda^{2}m\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+m\left\|y_{mh+j}-y_{mh+j}^{*}\right\|^{2},$
then
$\delta_{k}\leq$ $~{}\lambda^{2}m\omega_{\varepsilon}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+\frac{L_{gy}^{2}}{\mu_{g}^{2}}\tilde{\kappa}\left(j\right)\left(\frac{\sqrt{\tilde{\kappa}\left(j\right)}-1}{\sqrt{\tilde{\kappa}\left(j\right)}+1}\right)^{2j}\left\|v_{mh}-v_{mh}^{*}\right\|^{2}$
$+\frac{\tilde{\kappa}\left(j\right)}{\mu_{g}^{2}}\left(\frac{\mu_{g}L_{fy}+C_{fy}L_{gyy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\right)^{2}\left(\frac{\sqrt{\tilde{\kappa}\left(j\right)}-1}{\sqrt{\tilde{\kappa}\left(j\right)}+1}\right)^{2j}\left\|x_{mh+1}-x_{mh}\right\|^{2}$
$+\left(1+\sigma\right)^{j}\left(1-\theta\mu_{g}\right)^{j}\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}+m\right)\left\|y_{mh}-y_{mh}^{*}\right\|^{2}$
$+\left(1+\frac{1}{\sigma}\right)\left(1-\theta\mu_{g}\right)\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}+m\right)\left(\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{r=0}^{j}{\left(1+\sigma\right)^{r}\left(1-\theta\mu_{g}\right)^{r}\left\|x_{k-r}-x_{k-r-1}\right\|^{2}}.$ (68)
Then we can rearrange (H),
$\delta_{k}\leq(1-\gamma)\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+3m^{2}\lambda^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+\left\|{\widetilde{\nabla}\varphi(x_{mh})}\right\|^{2}\right),$ (69)
where we denote
$\omega_{\varphi}=\max$$\left\{\frac{\omega_{\varepsilon}}{m}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2},\frac{\tilde{\kappa}\iota}{m^{2}\mu_{g}^{2}}\left(\frac{\mu_{g}L_{fy}+C_{fy}L_{gyy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\right)^{2},\right.$
@@
Since
@@
$=$ $~{}3L_{gx}^{2}\delta_{k},$ (70)
it follows that
$\delta_{k}\leq$ $\left(1-\gamma\right)\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+3m^{2}\lambda^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+\left\|{\widetilde{\nabla}\varphi(x_{mh})}\right\|^{2}\right)$
$\leq$ $\left(1-\gamma\right)\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+6\lambda^{2}m^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)-\nabla\varphi(x_{mh+t})\right\|}^{2}+\left\|{\widetilde{\nabla}\varphi(x_{mh})-\nabla\varphi(x_{mh})}\right\|^{2}\right)$
$+6\lambda^{2}m^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+\left\|{{\nabla}\varphi(x_{mh})}\right\|^{2}\right)$
$\leq$ $\left(1-\gamma\right)\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+18\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\delta_{mh+t}}+\delta_{mh}\right)$
$+6\left(1-\gamma\right)\left(1-\gamma\right)^{-1}\lambda^{2}m^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|{\nabla}\varphi\left(x_{mh+t}\right)\right\|}^{2}+\left\|{{\nabla}\varphi(x_{mh})}\right\|^{2}\right).$ (71)
Based on (H), the following inequality can be checked by induction.
$\delta_{k}\leq\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+6\left(1-\gamma\right)^{-1}\lambda^{2}m^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}+\left\|{\nabla\varphi\left(x_{mh}\right)}\right\|^{2}\right),$
if $\lambda\leq\mathcal{O}\left(m^{-1}\iota^{m}\right)$, specifically,
$\lambda\leq\min\left\{\left(\frac{\gamma}{18\omega_{\varphi}}\right)^{1/2}\frac{1}{L_{gx}}m^{-3/2},\ \left(\frac{\gamma}{18\omega_{\varphi}}\right)^{1/2}\frac{1}{L_{gx}}m^{-1}\iota^{m-m_{0}}\right\}.$
Note that we can adjust the step size $\theta$ in (67) such that $\iota$ is closed to $1$, thus for simplicity, we still take $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$ We can bound $\delta_{k}$ recursively by
$\delta_{k}\leq$ $~{}\iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+6\left(1-\gamma\right)^{-1}\lambda^{2}m^{2}\omega_{\varphi}\sum_{t=0}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}$
$\leq$ $~{}\iota^{2\left(j-m_{0}\right)+h\left(m-m_{0}\right)}\delta_{0}+6\left(1-\gamma\right)^{-1}\lambda^{2}m^{2}\omega_{\varphi}\sum_{t=0}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}$
$+6\left(1-\gamma\right)^{-1}\lambda^{2}m^{2}\omega_{\varphi}\sum_{e=0}^{h-1}{\iota^{2\left(h-e\right)\left(m-m_{0}\right)}\sum_{t=0}^{m-1}{\left\|\nabla\varphi\left(x_{me+t}\right)\right\|^{2}}},$
where we are slightly abusing of notation that $\delta_{0}=\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{0}-y_{0}^{*}\right\|^{2}+\left\|v_{0}-v_{0}^{*}\right\|^{2}$ for analysis convenience, while we start from $(x_{1},y_{1},v_{1})$ in Algorithm 2 for the statement convenience in this paper. Substituting (H) into (H) completes the proof.
~$\left\|\widetilde{\nabla}\varphi\left(x_{k}\right)-\nabla\varphi\left(x_{k}\right)\right\|^{2}\leq$ $~{}3L_{gx}^{2}\delta_{k}$
$\leq$ $~{}3L_{gx}^{2}\iota^{2\left(j-m_{0}\right)+h\left(m-m_{0}\right)}\delta_{0}+18\left(1-\gamma\right)^{-1}\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\sum_{t=0}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}$
$+18\left(1-\gamma\right)^{-1}\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\sum_{e=0}^{h-1}{\iota^{2\left(h-e\right)\left(m-k\right)}\sum_{t=0}^{m-1}{\left\|\nabla\varphi\left(x_{me+t}\right)\right\|^{2}}}.$
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.9 3.10, and 3.11 hold. Within each epoch, we set the step size $\theta\sim\mathcal{O}(\frac{1}{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$, then the iterates $\{x_{k}\}$ generated by Algorithm 2 satisfy
$\frac{m}{K\left(m-m_{0}\right)}\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,\emph{{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}\leq\mathcal{O}\left(\frac{m\lambda^{-1}}{K\left(m-m_{0}\right)}\right),$
where $m_{0}\sim\mathcal{O}(\log m)$ is a constant and $m$ is the subspace dimension.
@@
According to Lemma E.2, a gradient descent step results in the decrease in the hyper-gradient $\varphi$:
$\varphi\left(x_{k+1}\right)\leq$ $~{}\varphi\left(x_{k}\right)+\left<\nabla\varphi\left(x_{k}\right),x_{k+1}-x_{k}\right>+\frac{L_{\varphi}}{2}\left\|x_{k+1}-x_{k}\right\|^{2}$
$\leq$ $~{}\varphi\left(x_{k}\right)-\left(\frac{\lambda}{2}-\lambda^{2}L_{\varphi}\right)\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}+\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)\left\|\nabla\varphi\left(x_{k}\right)-\widetilde{\nabla}\varphi\left(x_{k}\right)\right\|^{2}$
@@
$\leq$ $~{}\varphi\left(x_{k}\right)-\left(\frac{\lambda}{2}-\lambda^{2}L_{\varphi}\right)\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}+3\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)L_{gx}^{2}\iota^{2\left(j-m_{0}\right)+2h\left(m-m_{0}\right)}\delta_{0}$
@@
$+18\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)\left(1-\gamma\right)^{-1}\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\left(\sum_{t=0}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}+\sum_{e=0}^{h-1}{\iota^{2\left(h-e\right)\left(m-m_{0}\right)}\sum_{t=0}^{m-1}{\left\|\nabla\varphi\left(x_{me+t}\right)\right\|^{2}}}\right).$
Then, telescoping index from $k=0$ to $k=K=mH$,
$\left(\frac{\lambda}{2}-\lambda^{2}L_{\varphi}\right)\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}$
$\leq$ $~{}\varphi\left(x_{0}\right)-\varphi\left(x^{*}\right)+3\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)L_{gx}^{2}\delta_{0}\sum_{h=0}^{H-1}{\iota^{2h\left(m-m_{0}\right)}\sum_{j=m_{0}+1}^{m}{\iota^{2\left(j-m_{0}\right)}}}$
$+18\lambda^{2}\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)\left(1-\gamma\right)^{-1}L_{gx}^{2}m^{2}\omega_{\varphi}\left(m-m_{0}\right)\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}\sum_{h=0}^{H-1}{\iota^{2h\left(m-m_{0}\right)}}}$
$\leq$ $~{}\varphi\left(x_{0}\right)-\varphi\left(x^{*}\right)+3\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)L_{gx}^{2}\delta_{0}\frac{1-\iota^{2H\left(m-m_{0}\right)}}{1-\iota^{2}}$
$+18\lambda^{2}\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)\left(1-\gamma\right)^{-1}L_{gx}^{2}m^{2}\omega_{\varphi}\left(m-m_{0}\right)\frac{1-\iota^{2H\left(m-m_{0}\right)}}{1-\iota^{2}}\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}},$
where $x^{*}\in\operatorname*{arg\,min}_{x\in\mathbb{R}^{d_{x}}}\varphi(x)$. By selecting a suitable value for $\lambda$ such that
@@
$\lambda L_{\varphi}+18\lambda\left(\frac{\lambda}{2}+\lambda^{2}L_{\varphi}\right)\left(1-\gamma\right)^{-1}L_{gx}^{2}m^{2}\omega_{\varphi}\left(m-m_{0}\right)\frac{1-\iota^{2H\left(m-m_{0}\right)}}{1-\iota^{2}}\leq\frac{1}{4},$
we can derive
@@
$\sum_{\footnotesize\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}\leq\frac{4\left(\varphi\left(x_{0}\right)-\varphi\left(x^{*}\right)\right)}{\lambda}+\left(6+12\lambda L_{\varphi}\right)L_{gx}^{2}\delta_{0}\frac{1-\iota^{2H\left(m-m_{0}\right)}}{1-\iota^{2}},$
which completes the proof by dividing both sides by $\frac{m-m_{0}}{m}T$. ∎

### v2

@@ Appendix H Proof of The Main Theorem
Additionally, set the warm-up steps $m_{0}$ to satisfy
$\iota^{-2m_{0}}\geq\max\Bigg{\{}\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)^{-1}\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}+m\right),$
[…]
which means $m_{0}\sim\Omega\left(\log m\right)$. In this manner, adding $\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{k}-y_{k}^{*}\right\|^{2}$ on both sides of (H) and incorporating the estimation for $\epsilon_{j}^{h}$ (66) yield that
$\delta_{k}\leq$ $\ \iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+\left\|y_{mh}-y_{mh}^{*}\right\|^{2}+6m^{2}\lambda^{2}\omega_{\varphi}\left(\sum_{t=m_{0}}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+t}\right)\right\|^{2}}+\left\|\widetilde{\nabla}\varphi(x_{mh})\right\|^{2}\right)$
$\leq$ $\ \iota^{2\left(j-m_{0}\right)}\left(\delta_{mh}\right)+\iota^{2m}\|y_{m\left(h-1\right)}-y_{m\left(h-1\right)}^{*}\|^{2}$
$+6m^{2}\lambda^{2}\omega_{\varphi}\Big{(}\sum_{t=m_{0}}^{j-1}{\|\widetilde{\nabla}\varphi(x_{mh+t})\|^{2}}+\|\widetilde{\nabla}\varphi(x_{mh})\|^{2}$
$\quad\quad\quad\quad\quad+\sum_{t=m_{0}}^{m-1}{\|\widetilde{\nabla}\varphi(x_{m(h-1)+t})\|^{2}}+\|\widetilde{\nabla}\varphi(x_{m(h-1)})\|^{2}\Big{)},$ (68)
where we apply Lemma H.1 recursively to obtain the second inequality, and adopt the notation
$\omega_{\varphi}:=\max$$\left\{\frac{\omega_{\varepsilon}}{m}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2},\frac{\tilde{\kappa}\iota}{m^{2}\mu_{g}^{2}}\left(\frac{\mu_{g}L_{fy}+C_{fy}L_{gyy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\right)^{2},\right.$
@@
Consequently, expanding $\delta_{mh}$ in (68) in the same way derives (64). Regrading the upper-bound of the hyper-gradient estimation error, we have
@@
$=$ $~{}3L_{gx}^{2}\delta_{k}.$
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. Within each epoch, if we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$, and the step size for $x$ as zero in the first $m_{0}$ steps and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, the iterates $\{x_{k}\}$ generated by Algorithm 2 satisfy
$\frac{m}{K\left(m-m_{0}\right)}\sum_{\begin{subarray}{c}k=0,\\ (k\,\emph{{mod}}\,m)>m_{0}\end{subarray}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}=\mathcal{O}\left(\frac{m\lambda^{-1}}{K\left(m-m_{0}\right)}\right),$
where $m_{0}\sim\Omega(\log m)$ is a constant and $m$ is the subspace dimension.
@@
Consider the Lyapunov function $\mathcal{L}_{k}:=\varphi(x_{k})+\delta_{k}$. According to Lemma E.2, a gradient descent step leads to the decrease in the hyper-function:
$\varphi\left(x_{k+1}\right)-\varphi\left(x_{k}\right)\leq$ $\nabla\varphi\left(x_{k}\right),x_{k+1}-x_{k}\right>+\frac{L_{\varphi}}{2}\left\|x_{k+1}-x_{k}\right\|^{2}.$ (69)
Then, telescoping $\mathcal{L}_{k+1}-\mathcal{L}_{k}$ over the index set :=$\mathcal{I}:=\{k:\,0\leq k\leq K,\,(k\,{\texttt{mod}}\,m)>m_{0}\}$,
$\ \sum_{k\in\mathcal{I}}{\mathcal{L}_{k+1}-\mathcal{L}_{k}}$
$=$ $\ \sum_{k\in\mathcal{I}}{\varphi(x_{k+1})-\varphi(x_{k})}+\!\!\!\!\!\!\!\!\!\sum_{{\begin{subarray}{c}k=1,\\ ((k-1)\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}}^{K+1}{\!\!\!\!\!\!\!\!\!\delta_{k}}\ -\sum_{k\in\mathcal{I}}{\delta_{k}}$
@@
$\overset{(i)}{\leq}$ $\ \sum_{k\in\mathcal{I}}{\left<\nabla\varphi(x_{k}),x_{k+1}-x_{k}\right>+(\frac{L_{\varphi}}{2}+36m^{3}\omega_{\varphi})\|x_{k+1}-x_{k}\|^{2}}+\sum_{e=0}^{h-1}{2m\iota^{2(m-m_{0})}\delta_{me}}-\sum_{k\in\mathcal{I}}{\delta_{k}}$
@@
$\overset{(ii)}{\leq}$ $\ \sum_{k\in\mathcal{I}}-\left({\frac{\lambda}{2}-\lambda^{2}(L_{\varphi}+72m^{3}\omega_{\varphi})}\right)\|\nabla\varphi(x_{k})\|^{2}+\left({\frac{\lambda}{2}+\lambda^{2}(L_{\varphi}+72m^{3}\omega_{\varphi})}\right)\|\nabla\varphi(x_{k})-\widetilde{\nabla}\varphi(x_{k})\|^{2}$
$+\sum_{e=0}^{h-1}{2m\iota^{2(m-m_{0})}\delta_{me}}-\sum_{k\in\mathcal{I}}{\delta_{k}}$
@@
$\overset{(iii)}{\leq}$ $\ \sum_{k\in\mathcal{I}}{-\left({\frac{\lambda}{2}-\lambda^{2}(L_{\varphi}+72m^{3}\omega_{\varphi})}\right)\|\nabla\varphi(x_{k})\|^{2}}$
@@
$+3L_{gx}^{2}\left({\frac{\lambda}{2}+\lambda^{2}(L_{\varphi}+72m^{3}\omega_{\varphi})}\right)\sum_{k\in\mathcal{I}}{\delta_{k}}+\sum_{e=0}^{h-1}{2m\iota^{2(m-m_{0})}\delta_{me}}-\sum_{k\in\mathcal{I}}{\delta_{k}},$ (70)
where $(i)$ follows from inequalities (64) and (69); $(ii)$ results from the update rule $x_{k+1}=x_{k}-\lambda\widetilde{\nabla}\varphi(x_{k})$ and Young’s inequality; $(iii)$ comes from (63). Taking the coefficients of $\delta_{k}$ into account, we set the dimension parameters $(m,m_{0})$ satisfying $m\iota^{2\left(m-m_{0}\right)}<1/4$ and the step size $\lambda$ such that
$\lambda\leq\min\left\{\frac{1}{6L_{gx}^{2}},\ \frac{1}{\left(12\left(L_{\varphi}+72m^{3}\omega_{\varphi}\right)L_{gx}^{2}\right)^{1/2}},\ \frac{1}{4\left(L_{\varphi}+72m^{3}\omega_{\varphi}\right)}\right\}.$ (71)
In this way, we obtain the following result from (70),
$\mathcal{L}_{K+1}-\mathcal{L}_{0}=\sum_{{\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}}^{K}{\mathcal{L}_{k+1}-\mathcal{L}_{k}}\leq\sum_{{\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}}^{K}{-\left(\frac{\lambda}{2}-\lambda^{2}\left(L_{\varphi}+72m^{3}\omega_{\varphi}\right)\right)\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}.$
Rearrange the above inequality and denote $\varphi^{*}:=\min_{x\in\mathbb{R}^{d_{x}}}\varphi(x)$,
$\sum_{{\begin{subarray}{c}k=0,\\ (k\,{\texttt{mod}}\,m)>m_{0}\end{subarray}}}^{K}{\left\|\nabla\varphi\left(x_{k}\right)\right\|^{2}}\leq\frac{4\left(\varphi\left(x_{0}\right)-\varphi^{*}\right)}{\lambda}+\frac{4\delta_{0}}{\lambda},$
which completes the proof by dividing both sides by $\frac{m-m_{0}}{m}K$. ∎